# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Here we investigate analogy between quantum signal processing (QSP) and the adiabatic-impulse model (AIM) in order to implement the QSP algorithm with fast quantum logic gates. QSP is an algorithm that uses single-qubit dynamics to perform a polynomial function transformation. AIM effectively describes the evolution of a two-level quantum system under strong external driving field. We can map parameters from QSP to AIM to implement QSP-like evolution with non-adiabatic, high-amplitude external drives. By choosing AIM parameters that control non-adiabatic transition parameters (such as driving amplitude $A$, frequency $\omega$, and signal timing), one can achieve polynomial approximations and increase robustness in quantum circuits. The analogy presented here between QSP and AIM can be useful as a way to directly implement the QSP algorithm on quantum systems and obtain all the benefits from the fast Landau-Zener-Stückelberg-Majotana (LZSM) quantum logic gates.
@@ I Introduction
Quantum algorithms represent an innovative paradigm in computational science, using the principles of superposition, entanglement, and quantum interference to confront overwhelmingly time-consuming problems for classical approaches. Beyond the standard narratives of quantum speedup, quantum algorithms challenge assumptions in complexity theory by redefining the boundaries of efficiently solvable problems, such as Hamiltonian simulation, quantum search, quantum factoring, and quantum walks [1, 2, 3, 4, 5, 6, 7].
Quantum Signal Processing (QSP) is one such algorithm exploiting single-qubit dynamics with interference, enabling polynomial function transformations with $O(d)$ elementary unitary operations on the quantum subsystem, where $d$ is the degree of the polynomial function [8]. QSP algorithms have already undergone several improvements, leading to the development of new algorithms based on it, that enhance performance, address QSP problems [9] and can also be used as a part of more complex algorithms such as quantum search, Hamiltonian simulation, quantum phase estimation, etc. [5, 8]. QSP involves a combination of adjustable rotations around the $z$-axis, which produce phase differences between energy levels, and fixed $x$-axis rotations on the Bloch sphere, which change the energy levels occupation. These $x$- and $z$-axes rotations are applied one by one in a sequence. Such QSP sequences of pulses can be used as a component of a composite pulse approach that allows to increase robustness to driving signal noise [10].
We found that the combination of consecutive rotations around the $x$- and $z$-axes is also directly described by the so-called adiabatic-impulse model (AIM) [11], a model specifically developed for non-adiabatic single-qubit quantum dynamics, with particular emphasis on interferometry [12, 13]. It is a useful tool to investigate analytically the dynamics of quantum systems with a small energy gap [13, 14]. It combines adiabatic evolution, where the energy levels occupation is conserved but a phase difference is gained, and diabatic (impulse-type) transitions, during which the interference of the amplitudes of different states occurs. Thus, AIM results in a combination of rotations around the $x$- and $z$-axes, as in QSP. We find it beneficial to explore the analogy between the two approaches—QSP and AIM—and demonstrate that parameters obtained from AIM can be used to effectively implement QSP.
The rest of the paper is organized as follows. In Sec. II, we review the key notions of the work: QSP and AIM. In Sec. III, we describe the algorithms for converting parameters from QSP to AIM. In Sec. IV, we show the way to use two Landau-Zener-Stückelberg-Majorana (LZSM) transitions [12, 13, 15] to implement QSP with reduced duration. In addition, we show an analogy between double LZSM transitions and a Mach-Zehnder interferometer to implement a rotation operator around the $x$-axis by an arbitrary angle.
@@ II.1 Quantum Signal Processing
Quantum Signal Processing is a quantum algorithm initially designed for implementing polynomial transformations of matrix eigenvalues via phase modulation [6], originally developed for nuclear magnetic resonance applications [16]. It can also be used to enhance signal sensitivity to specific input parameters, making it a powerful tool for error mitigation, quantum optimization, and quantum machine learning tasks [17, 18]. For instance, the algorithm can be applied to amplify weak signals beyond the limits of repetition statistics and suppress noise without direct measurement [16].
The QSP algorithm involves using a sequence of unitary rotations of a qubit state on the Bloch sphere around the $x$- and $z$-axes. The operator responsible for rotations around the $x$-axis is denoted as the signal rotation operator $W$, and the operator responsible for rotation around the $z$-axis is called the signal processing operator $S$
@@
where $X$ and $Z$ stand for the Pauli matrices $\sigma_{x}$ and $\sigma_{z}$, respectively. Here the parameter $a$ is defined as
@@
the operator $U$, which represents the evolution in the QSP algorithm, is defined as the sequential application of the signal processing operator $S$ and the signal rotation operator $W$:
@@
Therefore, the operators $W$ and $S$ can be written in the following way using rotation operators
@@
Quantum rotation gates $R_{x}$ and $R_{z}$ describe rotations of single-qubit states on the Bloch sphere by angles $\alpha$ and $\beta$ around the $x$- and $z$-axes, respectively. These gates are defined both in terms of matrix exponentials and explicit matrix representations
@@
The equation (4) can be written in terms of rotation operators $R_{x}$ and $R_{z}$
$U_{\boldsymbol{\Phi},a}=R_{z}(-2\phi_{0})\prod_{k=1}^{d}R_{x}(-\theta)R_{z}(-2\phi_{d}).$ (7)
One of the most remarkable properties of the QSP framework is its ability to construct quantum circuits that implement a desired polynomial transformation on the eigenvalues of an operator. The inverse of this principle is also true: given a target polynomial $M(a)$, there exists a corresponding set of QSP phase angles $\boldsymbol{\Phi}$ such that the transformation $M(a)$ can be exactly represented as $\langle 0|U_{\boldsymbol{\Phi},a}|0\rangle$. This principle is formalized in Theorem 1 (Quantum Signal Processing) of Ref. [8], which states that the QSP sequence $U_{\boldsymbol{\Phi},a}$ generates a matrix elements, expressible via polynomial function of $a$:
@@
For this construction to hold, the polynomials $M(a)$ and $Q(a)$ must satisfy several conditions. The degree of the polynomial $M(a)$ must be at most $d$ and the degree of $Q(a)$ must be at most $(d-1)$. The polynomials must also exhibit parity relationships with the degree $d$, and a normalization condition must ensure the unitarity of the resulting transformation. Moreover, for any polynomials $M(a)$ and $Q(a)$ satisfying the feasibility conditions, such a sequence of phase angles $\boldsymbol{\Phi}$ always exists.
Figure 1: Occupation probability $P_{-}$ after the QSP sequence applied to the qubit with ground state initial conditions. Chebyshev polynomials with Eqs. (9b,9c) are shown in panel (a). The BB1 sequence, corresponding to the polynomial in Eq. (12), is shown in panel (b) (solid curves). Orange dots correspond to the probability obtained by performing the respective QSP sequences on an IBM-Q quantum processing unit (QPU). The scheme for the BB1 sequence is shown in terms of QSP operators, where green boxes correspond to the signal processing operator $S(\phi)$ and blue boxes correspond to the signal rotation operator $W(a)$.
As an example, we now consider how Chebyshev polynomials $T_{d}(a)$ [19] naturally arise within the framework of QSP [8] and enable polynomial transformations. For trivial phase sequences (i.e., all phases in $\boldsymbol{\Phi}$ are zero), the resulting polynomials $M(a)$ align exactly with $T_{d}(a)$:
@@
These results demonstrate that QSP can directly generate the Chebyshev polynomials, which are defined recursively as
$T_{0}(a)=1,\hskip 10.00002ptT_{1}(a)=a;$ (10a)
$T_{d}(a)=2aT_{d-1}(a)-T_{d-2}(a)=\cos(d\arccos a).$ (10b)
Another important QSP sequence is BB1 (Broadband 1) [16]. It is a robust composite pulse sequence used in NMR to correct amplitude errors in RF pulses. It is particularly useful for achieving high-fidelity spin manipulations, even when the RF field strength is not perfectly calibrated. To perform the BB1 sequence, we use
@@
where $\text{where }\eta=\frac{1}{2}\cos^{-1}\left(-1/4\right),$
which results in the polynomial:
@@
The BB1 sequence describes a qubit that remains in the excited state in the broad region $\theta\approx(-2/3\pi,2/3\pi)$, and quickly becomes flipped out of it.
In Fig. 1 we illustrate the resulting probability after applying the QSP sequence with the trivial sequences for the Chebyshev polynomials in panel (a) and BB1 phase sequence in panel (b). A comparison is made between these results and those obtained from running the corresponding QSP sequences on a quantum computer via IBM-Q cloud quantum computing on the QPU based on the IBM Eagle r3 architecture [20]. We obtained good agreement for this simple single-qubit algorithm despite dissipative and noisy environment of the real qubits in the IBM-Q processor.
@@
Later, in Sec. III, we demonstrate how closely this algorithm resembles the adiabatic-impulse model.
@@ II.2 Adiabatic-impulse model
The adiabatic-impulse model [11, 21, 22, 23, 24, 25] describes the evolution of a system driven by a periodic signal with a linear transition through anti-crossing point (point of the minimal energy gap). This model works well with systems with small energy gap and can be used as a basis for quantum logic gates [26, 27, 14, 28]. The evolution is divided into two regimes: the adiabatic regime, where the system closely follows its adiabatic states, and the diabatic impulse-type regime, where a transition of occupation probability between two energy levels occurs, accompanied by a phase gain between different energy level occupations.
@@
where $\Delta$ is the minimal energy gap between the two energy levels and $\varepsilon(t)=-A\cos\omega t$ is the driving signal with frequency $\omega$ and large amplitude $A>\Delta$. The evolution is separated into adiabatic evolution, described by the matrix $U$, and transition (impulse-type) evolution described, by the matrix $N$:
@@
$N=\begin{pmatrix}\mathcal{R}e^{-i\phi_{\text{S}}}&-\mathcal{T}\\ \mathcal{T}&\mathcal{R}e^{i\phi_{\text{S}}}\end{pmatrix}=R_{z}(\phi_{\text{S}})R_{x}(\theta)R_{z}(\phi_{\text{S}}),$ (14b)
$\phi_{\text{S}}=\dfrac{\pi}{4}+\delta(\ln{\delta}-1)+\arg{\left[\Gamma(1-i\delta)\right]},\hskip 10.00002pt\delta=\dfrac{\Delta^{2}}{4\hbar v},$
$\mathcal{T}=\sqrt{\mathcal{P}},\hskip 10.00002pt\mathcal{R}=\sqrt{1-\mathcal{P}},\ \mathcal{P}=\exp{\left[\dfrac{-\pi\Delta^{2}}{2A\hbar\omega}\right]}.$ (14c)
where $\phi_{\text{S}}$ is the Stokes phase, $v=\frac{d\varepsilon(t)}{dt}|_{\varepsilon=0}=A\omega$ is the speed of the anti-crossing passage (first derivative of the driving signal in the anti-crossing point), $\mathcal{T}$ is the transmission coefficient that is equivalent to $a$ in the QSP algorithm, so that the polynomial base in QSP is the transmission coefficient. Moreover $\mathcal{R}$ is the reflection coefficient, $\mathcal{P}$ is the LZSM excitation probability of the qubit with a single transition from the ground state [29, 30, 31, 32, 33, 34, 35], and $\delta$ is the adiabaticity parameter. Note that the transition matrix $N$ depends on the direction of the crossing of the anti-crossing region and in the case of inverse crossing, the matrix $N^{\text{inv}}$ includes an additional phase $\pi$ along with the Stokes phase [13, 24], or in other words it equals to the transposed transition matrix $N^{\text{inv}}=N^{\top}$. The phase, accumulated during the adiabatic evolution of the system, is:
[…]
@@
According to the AIM, the evolution of the system is adiabatic everywhere except at the LZSM transition point (or anti-crossing point), while the transition itself is considered instantaneous in the AIM, and the evolution during this time is described as diabatic. In this case, the complete evolution of the system during a single LZSM transition, from the initial state $\psi_{\text{in}}$ (index 1) to the final state $\psi_{\text{fin}}$ (index 2) consists of adiabatic evolution before the anti-crossing point, transition evolution at the anti-crossing point, and adiabatic evolution after the anti-crossing point:
@@
where $\zeta_{1}$ and $\zeta_{2}$ are the phase gains before and after the anti-crossing point, respectively.
The complete evolution matrix for a single transition in terms of Bloch sphere rotation operators is
$U_{\text{LZSM}_{1\rightarrow{}2}}=R_{z}(2\zeta_{2}+\phi_{\text{S}})R_{x}(\theta)R_{z}(2\zeta_{1}+\phi_{\text{S}}).$ (17)

### v3

@@ Abstract
We investigate an analogy between quantum signal processing (QSP) and the adiabatic-impulse model (AIM) in order to implement the QSP algorithm with fast quantum logic gates. QSP is an algorithmic technique that uses single-qubit dynamics to perform a polynomial function transformation. The adiabatic-impulse model effectively describes the evolution of a two-level quantum system under a strong external driving field. We can map parameters from QSP to the AIM to implement a QSP-like evolution with non-adiabatic, high-amplitude external drives. By choosing the AIM driving parameters that control the non-adiabatic transition parameters (such as driving amplitude $A$, frequency $\omega$, and signal timing), one can achieve polynomial approximations and increase robustness in quantum circuits. The analogy between QSP and AIM presented here can be useful as a way to directly implement the QSP algorithmic technique on quantum systems. And show the benefits from fast Landau-Zener-Stückelberg-Majorana (LZSM) quantum logic gates in comparison with usual resonance driving gates in IBM-Q quantum processors.
@@ I Introduction
Quantum algorithms represent an innovative paradigm in computational science, using the principles of superposition, entanglement, and quantum interference to confront overwhelmingly time-consuming problems for classical approaches [1, 2, 3, 4, 5, 6, 7, 8]. Quantum algorithms have been reinterpreted within more general frameworks that emphasize common structural principles. A notable example is the perspective framework, developed in Ref. [9], named “grand unification of quantum algorithms”, which shows that a wide range of fundamental algorithms can be understood as instances of singular value transformations. At the core of this view lies Quantum Signal Processing (QSP), a method exploiting single-qubit dynamics with interference, enabling polynomial function transformations with $O(d)$ elementary unitary operations on the quantum subsystem, where $d$ is the degree of the polynomial function. The role of QSP as a unifying primitive has motivated its integration into increasingly sophisticated algorithmic designs [10, 5, 11], where it functions both as a direct tool and as a foundation for broader constructions such as Quantum Singular Value Transformation (QSVT).
QSP involves a combination of adjustable rotations of a qubit state around the $z$-axis, which produces phase differences between energy levels, and fixed $x$-axis rotations on the Bloch sphere, which change the energy levels occupation. These $x$- and $z$-axes rotations are applied one by one in a sequence. Such QSP sequences of pulses can be used in a composite-pulse approach that allows to increase robustness to fluctuations in the driving signal [12].
We found that the combination of consecutive rotations around the $x$- and $z$-axes is also directly used in the so-called adiabatic-impulse model (AIM) [13], a model specifically developed for non-adiabatic quantum dynamics of qubits, with particular emphasis on interferometry [14, 15, 16, 17]. It is a useful tool to investigate analytically the dynamics of quantum systems with a small energy gap [17, 18, 19]. It combines adiabatic evolution, where the energy levels occupation is conserved but a phase difference is gained, and diabatic (impulse-type) transitions, during which the interference of the amplitudes of different states occurs.
We find it beneficial to explore the analogy between the two approaches—QSP and AIM—and demonstrate that the AIM can be used to effectively implement QSP. This is possible by utilizing alternative fast non-adiabatic gates [20].
The rest of the paper is organized as follows. In Sec. II, we review the key notions of this work: QSP and AIM. In Sec. III, we describe the algorithms for converting parameters from QSP to AIM. In Sec. IV, we show the way to use two consecutive Landau-Zener-Stückelberg-Majorana (LZSM) transitions [14, 17, 21] to implement QSP faster. In addition, we show an analogy between double LZSM transition and a Mach-Zehnder interferometer to implement a rotation operator around the $x$-axis by an arbitrary angle.
@@ II.1 Quantum Signal Processing
Quantum Signal Processing is a quantum algorithmic technique initially designed for implementing polynomial transformations of matrix eigenvalues via phase modulation, originally inspired by composite-pulse techniques from nuclear magnetic resonance [22]. The possibility to implement such polynomial transformations is particularly important in the context of applications like Hamiltonian simulation, as we will demonstrate later. It can also be used to enhance signal sensitivity to specific input parameters, making it a powerful tool for error mitigation, quantum optimization, and quantum machine learning tasks [23, 24]. For instance, the technique can be applied to amplify weak signals beyond the limits of repetition statistics and suppress noise without direct measurement [22].
The QSP algorithm involves using a sequence of unitary rotations of a qubit state on the Bloch sphere around the $x$- and $z$-axes. The operator responsible for rotations around the $x$-axis is denoted as the signal operator $W$, and the operator responsible for rotation around the $z$-axis is called the signal processing operator $S$
@@
where $X$ and $Z$ stand for the Pauli matrices $\sigma_{x}$ and $\sigma_{z}$, respectively. Here, the parameter $a$ is defined as
@@
the operator $U$, which represents the evolution in the QSP technique, is defined as the sequential application of the signal processing operator $S$ and the signal operator $W$:
@@
The operators $W$ and $S$ can be written in the following way using rotation operators
@@
The quantum rotation gates $R_{x}$ and $R_{z}$ describe rotations of single-qubit states on the Bloch sphere by angles $\alpha$ and $\beta$ around the $x$- and $z$-axes, respectively. These gates are defined both in terms of matrix exponentials and explicit matrix representations
@@
Consequently, the equation (4) can be written in terms of rotation operators $R_{x}$ and $R_{z}$
$U_{\boldsymbol{\Phi},a}=R_{z}(-2\phi_{0})\prod_{k=1}^{d}R_{x}(-\theta)R_{z}(-2\phi_{k}).$ (7)
One of the most remarkable properties of the QSP framework is its ability to construct quantum circuits that implement a desired polynomial transformation on the eigenvalues of an operator.
The inverse of this principle is also true: given a target polynomial $M(a)$, there always exists a corresponding set of QSP phase angles $\boldsymbol{\Phi}$ such that the transformation $M(a)$ can be exactly represented as $\langle 0|U_{\boldsymbol{\Phi},a}|0\rangle$. This principle is formalized in Theorem 1 (Quantum Signal Processing) of Ref. [9], which states that the QSP sequence $U_{\boldsymbol{\Phi},a}$ generates a matrix elements, expressible via a polynomial function of $a$:
@@
For this construction to hold, the desired $M(a)$ and complementary $Q(a)$ polynomials must satisfy several conditions. The degree of the polynomial $M(a)$ must be at most $d$ and the degree of the polynomial $Q(a)$ must be at most (d1)(d-1). These polynomials must also exhibit parity relationships with the degree $d$, and a normalization condition must ensure the unitarity of the resulting transformation.
Figure 1: Occupation probability $P_{-}$ after the QSP sequence is applied to the qubit with ground state initial conditions. Chebyshev polynomials with Eqs. (10b,10c) are shown in panel (a). The BB1 sequence, corresponding to the polynomial in Eq. (12), is shown in panel (b) (solid curve). The dots in (a,b) correspond to the probability obtained by performing the respective QSP sequences on an IBM-Q quantum processing unit (QPU). The scheme for the BB1 sequence is shown (b) in terms of QSP operators, where the green boxes correspond to the signal processing operator $S(\phi)$ and blue boxes correspond to the signal operator $W(a)$.
As an example, we now consider the Chebyshev polynomials $T_{d}(a)$ [25], which are defined as follows
$T_{0}(a)=1,\hskip 10.00002ptT_{1}(a)=a;$ (9a)
$T_{d}(a)=2aT_{d-1}(a)-T_{d-2}(a)=\cos(d\arccos a),$ (9b)
and how they naturally arise within the framework of QSP [9]. Such polynomials play a central role in Hamiltonian simulation, since time-evolution operators of the form $\exp{(-iHt)}$ are often approximated through Chebyshev polynomial series. For trivial phase sequences (i.e., all phases in $\boldsymbol{\Phi}$ equal zero), the resulting polynomials $M(a)$ align exactly with $T_{d}(a)$:
@@
Another important QSP sequence is BB1 (Broadband 1) [22]. It is a robust composite pulse sequence used in NMR to correct amplitude errors in RF pulses. This is particularly useful for achieving high-fidelity spin manipulations, even when the RF field strength is not perfectly calibrated. Following Ref. [9] we consider BB1 with a phase sequence
@@
$\text{where }\eta=\frac{1}{2}\cos^{-1}\left(-\frac{1}{4}\right).$
Via the QSP technique, Eq. (4), one can obtain the polynomial form from the phase sequence in Eq. (11)
@@
The BB1 sequence in Eq. (11) describes a qubit that remains in the excited state in the broad region $\theta\approx(-2/3\pi,2/3\pi)$ and quickly becomes flipped out of it.
In Fig. 1 we illustrate the resulting probability after applying the QSP sequence with the trivial phases for the Chebyshev polynomials in panel (a) and BB1 phase sequence in panel (b). A comparison is made between these results (shown with solid curves) and those obtained from running the corresponding QSP sequences on an IBM-Q cloud quantum computing on the QPU based on the IBM Eagle r3 architecture [26], shown by color dots. We obtained good agreement for this single-qubit technique despite the dissipative and noisy environment of real qubits in the IBM-Q processor.
@@
Later, in Sec. III, we demonstrate how closely this technique resembles the adiabatic-impulse model.
@@ II.2 Adiabatic-impulse model
The adiabatic-impulse model [13, 27, 28, 29, 30, 31] describes the evolution of a system driven by a periodic signal with a linear transition through an anti-crossing point (point of the minimal energy gap). This model works well with systems with a small energy gap and can be used as a basis for quantum logic gates [20, 32, 18, 33, 34]. The evolution is divided into two regimes: the adiabatic regime, where the system closely follows its adiabatic states, and the diabatic impulse-type regime, where a transition of occupation probability between two energy levels occurs, accompanied by a phase gain between different energy level occupations.
@@
where $\Delta$ is the minimal energy gap between the two energy levels and $\varepsilon(t)=-A\cos\omega t$ is the driving signal with frequency $\omega$ and large amplitude $A>\Delta$. The evolution is separated into adiabatic evolution, described by the matrix $U$, and transition (impulse-type) evolution described by the matrix $N$:
@@
$N=\begin{pmatrix}\mathcal{R}\exp{(-i\phi_{\text{S}})}&-\mathcal{T}\\ \mathcal{T}&\mathcal{R}\exp{(i\phi_{\text{S}})}\ \end{pmatrix}$ (14b)
$=R_{z}(\phi_{\text{S}})R_{x}(\theta)R_{z}(\phi_{\text{S}}),$ (14c)
$\mathcal{T}=\sqrt{\mathcal{P}},\hskip 10.00002pt\mathcal{R}=\sqrt{1-\mathcal{P}},$ (14d)
$\phi_{\text{S}}=\dfrac{\pi}{4}+\delta(\ln{\delta}-1)+\arg{\left[\Gamma(1-i\delta)\right]},$ (14e)
$\mathcal{P}=\exp{\left[-2\pi\delta\right]},\hskip 10.00002pt\delta=\dfrac{\Delta^{2}}{4\hbar v},\hskip 10.00002pt$ (14f)
$\left.v=\frac{d\varepsilon(t)}{dt}\right|_{\varepsilon=0}=A\omega,$ (14g)
where $\phi_{\text{S}}$ is the Stokes phase, $v$ is the speed of the anti-crossing passage, $\Gamma$ is the gamma function, $\mathcal{T}$ is the transmission coefficient that is equivalent to a polynomial base $a$ in the QSP algorithmic technique. Moreover, $\mathcal{R}$ is the reflection coefficient, $\mathcal{P}$ is the LZSM excitation probability of the qubit with a single transition from the ground state [35, 36, 37, 38, 39, 40, 41], and $\delta$ is the adiabaticity parameter. Note that the transition matrix $N$ depends on the direction of the crossing of the anti-crossing region and in the case of inverse crossing, the matrix $N^{\text{inv}}$ includes an additional phase $\pi$ along with the Stokes phase [17, 30]; or in other words, it equals the transposed transition matrix $N^{\text{inv}}=N^{\top}$. The phase, accumulated during the adiabatic evolution of the system, is:
$\zeta(t_{i},t_{j})$ $=\dfrac{1}{2\hbar}\int^{t_{j}}_{t_{i}}\Delta E(t)dt,$ (15a)
@@
According to the AIM, the evolution of the system is adiabatic everywhere except at the LZSM transition point (or anti-crossing point), while the transition itself is considered instantaneous in the AIM, and the evolution during this time is described as diabatic. In this case, the complete evolution of the system during a single LZSM transition, from the initial state $\psi_{\text{in}}$ to the final state $\psi_{\text{fin}}$ consists of adiabatic evolution before the anti-crossing point, transition evolution at the anti-crossing point, and adiabatic evolution after the anti-crossing point:
@@
where $\zeta_{1}$ and $\zeta_{2}$ are the phase gains before and after the anti-crossing point, respectively. The complete evolution matrix for a single transition in terms of Bloch sphere rotation operators is
$U_{\text{LZSM}}=R_{z}(2\zeta_{2}+\phi_{\text{S}})R_{x}(\theta)R_{z}(2\zeta_{1}+\phi_{\text{S}}).$ (17)